the length of a rectangular field is 75yd.This is 3yd more than twice the width.how wide is the field??
Answer by checkley71(8403)

L=75
L=2W+3
2W+3=75
2X=75-3
2W=72
W=72/2
W=36 IS THE WIDTH.
PROOF:
75=2*36+3
75=72+3
75=75